# The NEID Earth Twin Survey. III. Survey Performance After Three Years on Sky

### v1

@@ 1 Introduction
The NEID Earth Twin Survey (NETS; Gupta et al., 2021) is a dedicated search for low-mass, long-period exoplanets with NEID (Schwab et al., 2016), an extreme precision radial velocity (EPRV) spectrograph on the WIYN 3.5 m telescope at Kitt Peak National Observatory.111The WIYN Observatory is a joint facility of the NSF’s National Optical-Infrared Astronomy Research Laboratory, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, and Princeton University. Using time allocated through a guaranteed time observations (GTO) program, supplemented in part by Penn State institutional access to NEID, we have been monitoring a carefully selected sample of 41 bright, RV-quiet stars since the commencement of this survey in September 2021. Data from the first few years of the survey have led to the confirmation of a low-mass planet (Gupta et al., 2025), the precise characterization of long-period stellar and planetary orbits (Giovinazzi et al., 2025), and the refinement of stellar activity mitigation methods (Burrows et al., 2024; Siegel et al., 2024; Gilbertson et al., 2024).
@@ 2.1 NETS Observing Strategy
NEID is a fiber-fed (Kanodia et al., 2023), environmentally-stabilized (Robertson et al., 2019) optical to near-infrared spectrograph. The instrument was designed to exceed an internal precision requirement of 27 cm s-1 based on a comprehensive, bottom-up error budget presented by Halverson et al. (2016). Tests performed during instrument commissioning and data from the NEID solar feed (Lin et al., 2022; Ford et al., 2024) demonstrate an achieved on-sky precision of cm s-1. In addition, thanks to its broad wavelength coverage of more than 380 – 930 nm, NEID observations cover a number of spectral features that are known to trace stellar magnetic activity, including the Ca II H&K lines, H, and the Ca II infrared triplet.
@@ 2.1.1 Nightly Observing Strategy
The target single-measurement precision for NETS observations follows that described by Gupta et al. (2021), which we summarize here. We first calculate how long it will take to reach a combined cm s-1. The photon noise precision, , is estimated for median observing conditions using the NEID exposure time calculator, which interpolates over a grid of exposure time, stellar effective temperature (), and apparent -mag generated following Bouchy et al. (2001). For the photon noise calculation, the projected rotational velocity, , was held fixed at 2 km s-1. The residual p-mode oscillation amplitude, , is estimated using the model outlined by Chaplin et al. (2019). We then compare this to the period corresponding to the peak of the oscillation power envelope, , and take the larger of these two values to be the nominal exposure time for each star. To avoid incurring additional uncertainty due to variations in the charge transfer inefficiency (CTI) of the NEID detector (Blake et al., 2017), individual exposures are set to trigger at a fixed signal-to-noise ratio (S/N) corresponding to this nominal exposure time. For the brightest stars in our sample, for which the calculated S/N exceeds the detector non-linearity threshold at any wavelength (60% well depth; S/N per 1-D extracted pixel), observations are split into multiple consecutive exposures with a total baseline (including inter-exposure readout) equal to the nominal exposure time. We also note that for all NETS observations, we simultaneously feed light from the etalon (Kreider et al., 2025) through the calibration fiber, except when the etalon is offline for maintenance. These measurements can be used to improve sampling of the nightly instrument drift.
@@
The 30 cm s-1 precision threshold was selected by Gupta et al. (2021) because it is on par with the internal NEID precision requirement and it can be achieved with reasonably short exposures ( minutes) for most NETS targets. For a subset of these stars, however, we expect to reach 30 cm s-1 with exposure sequences shorter than the -minute overheads associated with each NEID observation. To improve the duty cycle and increase the precision of observations for some of the brightest stars, we double the number of exposures per visit for the stars HD 4614A, HD 10476, HD 26965, HD 115617, HD 185144, and HD 201091, and we triple the number of exposures per visit for HD 10700. In each of these cases, the total nominal visit duration is still less than 10 minutes.
@@
While NEID commissioning was ongoing and prior to initiating observations for the full NETS sample in September 2021, we observed a limited set of stars from January 2021 through August 2021 to explore the precision capabilities of the spectrograph and the practicality of our intended nightly observing strategy within the constraints of the NEID Queue.222For details about the NEID Queue, see https://wiyn-queuemaster.kpno.noirlab.edu/manual/ For these precursor observations, we aimed to schedule two visits per night, separated by a minimum of two hours. This sampling was intended to mitigate the impact of correlated noise contributions from granulation and supergranulation (Dumusque et al., 2011), as these processes have characteristic timescales of hours to days. While these observations were executed as intended in most cases, we were sometimes left with unpaired visits as a result of changing observing conditions and unforeseen technical issues. In the interest of maintaining a more consistent observing strategy across all nights, we elected to proceed with just a single visit per night once the full survey began.
@@ 2.1.2 Seasonal Observing Strategy
For each star, we schedule P1 visits in the middle of each season at high cadence, set to execute as frequently as once every two nights. These observations, which typically span 1–2 months, allow us to densely sample the RV time series on timescales comparable to the rotation periods of our target stars. Rotationally-modulated stellar activity signals are expected to complicate exoplanet detection for even the quietest stars (Crass et al., 2021), but simultaneous measurement of RVs and activity indicators can be a powerful tool for disentangling stellar and planetary signals at the sub-m s-1 level. However, this requires that our data sets are sufficiently well sampled so that we can measure and model activity variations on the relevant timescales (Rajpaul et al., 2015; Haywood et al., 2018; Hall et al., 2018; Lanza et al., 2019; Yu et al., 2024; Gupta & Bedell, 2024). On their own, these densely-sampled observations would lead to poor phase sampling for longer period exoplanet signals, which can degrade the accuracy and precision of the extracted signal amplitude (Burt et al., 2018). To improve our phase sampling, we use our P2 allocation to schedule low cadence observations that span each star’s observing season. Finally, to reduce annual aliasing effects (which can be a significant obstacle to the detection of planet with periods close to one year), we schedule sets of 2–3 P0 visits as each star is rising or setting to bracket the observing season. By making use of the highest priority time for these observations, we ensure that the stars are observed within a few weeks of each end of the observing season when they are observable for no more than a couple hours each night. This counteracts the tendency for targets to drop out of the schedule during these limited observability windows. An idealized version of the seasonal observing schedule for a typical star is shown in the top panel of Figure 1. The schedule is supplemented by additional P3 and P4 time provided by Penn State; these observations follow the same nightly observing strategy, but no constraints are imposed on the observing cadence at these low priorities. We allow each star to be observed no more than once per night across all priority levels.
@@ 2.2 Current Status of NETS Observations
NETS observations formally commenced in September 2021, at the start of the 2021B observing semester. We show the achieved NETS observing schedule through August 2024 in Figure 1. In addition to the observations that follow the seasonal and nightly strategies outlined above, we also include supplemental data taken outside of this algorithmic scope and data from the NEID standard star program. The supplemental data adhere to the nightly NETS observing strategy (Section 2.1.1) and these include observations at lower priorities (P3 and P4) as well as observations scheduled on an ad hoc basis to monitor interesting signals or to compensate for weather losses and other scheduling setbacks. The target list for the NEID standard star program includes the following NETS targets: HD 4628, HD 9407, HD 10700, HD 127334, HD 185144, and HD 221354. The standard star program uses the same S/N trigger as NETS for all but HD 10700, for which a slightly lower threshold is used. For the analysis in this work, we combine data from both programs, as the RV uncertainty introduced by CTI variations is not expected to exceed 10 cm s-1 for observations at high S/N (Blackman et al., 2020). For more detailed analyses, one could consider separating the NETS and standard star data for HD 10700.
@@ 2.3 Data Products
This work uses NETS data as processed by version 1.4 of the NEID Data Reduction Pipeline (DRP)444Pipeline documentation can be found at https://neid.ipac.caltech.edu/docs/NEID-DRP/, which uses the cross-correlation function (CCF; Baranne et al., 1996) method with spectral type-dependent masks to calculate RVs. We retrieved the processed spectra from the NEID data archive.555https://neid.ipac.caltech.edu Data collected during Science Run 1 (2020 October 26 – 2022 June 16) and Science Run 2 (2022 October 18 – 2024 August 19) are treated as distinct RV eras, as the thermal cycling of the spectrograph in response to the Contreras fire is expected to have introduced an offset in the instrumental RV zero point666https://neid.ipac.caltech.edu/docs/NEID-DRP/rveras.html. We have also identified an apparent RV zero point offset in the middle of Science Run 1, the physical cause of which is yet unknown, so we consider data taken from 2020 October 26 – 2021 August 31 as a separate “Run 0.5”. We refer to these as Run 0.5, Run 1, and Run 2 in this work, and the offsets are described in detail in Section 3.3.
In addition to precise RVs, the DRP provides useful diagnostics such as the CCF full width at half maximum (FWHM) and bisector inverse slope (BIS) as well as telluric-corrected measurements of various indicators of stellar activity levels. These are calculated as the ratio of the flux in the specified activity-sensitive line core relative to a nearby continuum region, as described in the DRP documentation. Among these indicators is the Ca II H&K emission, though the values provided by the NEID DRP have not been calibrated to the standard Mount Wilson (MW) scale for the commonly used SHK metric (Duncan et al., 1991, also called the S-index or S-value). We follow the approach of Lovis et al. (2011) and Gomes da Silva et al. (2021) and compute the scaled SHK,MW values from the native DRP Ca II H&K activity indicator by performing a linear fit between the average Ca II H&K values from the NEID DRP and the mean S-values from Duncan et al. (1991) for the 27 FGK stars that appear in both samples. We find a conversion S Ca II H&K, and we expect this calibration to be valid for stars with spectral types and activity levels comparable to the input sample (FGK stars with Ca II H&K). We report both the DRP Ca II H&K values as well as the scaled SHK,MW for each star, with the exception of HD 95735, the sole M-dwarf on the NETS target list. The time series and generalized Lomb-Scargle (GLS; Zechmeister & Kürster, 2009, 2018) periodograms for several activity-sensitive lines are shown in Figure 2 for a typical NETS target, and in the accompanying figure set for the remaining target stars. False alarm probabilities (FAP) are calculated following Zechmeister & Kürster (2009), accounting for the number of independently sampled frequencies.
@@ 3.2 Stellar companions
Several of the stars in our sample are members of binary (or higher multiplicity) systems. To determine which of these stellar companions are expected to induce measurable RV signals, we use binary_mc (Giovinazzi, 2025) to simulate possible orbits based on the positions of each star as given by the Gaia DR3 catalog (Gaia Collaboration et al., 2023). We calculate the maximum RV slope for each of 10 million random orbit draws for each system, and we consider the slope to be measurable in systems for which at least 10% of orbits have maximum slopes m s-1 yr-1. Four of these, HD 4614 A, HD 24496 A, HD 68017, and HD 201091 (61 Cyg A), were analyzed in a study of NETS stars with high astrometric accelerations by Giovinazzi et al. (2025) which combined astrometric measurements, archival RVs, and Run 1 NETS data to refine the binary orbits. We adopt and subtract the published orbits from the observed RVs for each of these stars. The orbital parameters are given in Table 2. An additional three stars, HD 179957, HD 186408, and HD 116442 have measurable RV signals due to their stellar companions but did not meet the criteria for thorough characterization in Giovinazzi et al. (2025). For these stars, we fit and subtract separate linear trends from NEID time series for each run.
Other stars in our sample with stellar-mass companions include HD 86728, which has a low-mass M-dwarf binary (Gizis et al., 2000; Hirsch et al., 2021) and HD 26965, which is part of a triple star system in which the orbit of the B and C system is well characterized (Mason et al., 2017) but that of the BC pair about the primary star is not. For these systems, we neither expect nor observe measurable RV signals as a consequence of some combination of long orbital periods, low companion masses, and highly inclined orbits.
@@ 3.3 RV Zero Point Offset
The extraordinary stability of EPRV spectrographs such as NEID is achieved through careful environmental control with extremely strict tolerances (e.g., Stefansson et al., 2016; Robertson et al., 2019; Crass et al., 2022). As a consequence, changes to the reference wavelength scale and average line profile resulting from even minor changes to the spectrograph or its calibration sources may introduce a measurable RV zero point offset. Our analysis of the line profiles of NETS observations reveals that two such changes occurred during the first three years of the survey. In Figure 3, we show the residual CCFs for all spectra of the star HD 185144, which was observed as part of both the NEID standard star program and NETS. The residual CCFs are calculated as the difference the CCF of each observation and the average CCF over all observations. There are two distinct changes to the average line profile; the first occurs between 2021 August 7 and 2021 September 17, and the second occurs between 2022 June 10 and 2022 December 9. The latter change corresponds to the break between Run 1 to Run 2 and is explained by the derivation of a new master wavelength solution for the Thorium-Argon (ThAr) calibration lamp and by the thermal cycling of the instrument in response to the Contreras fire. The earlier break corresponds to dates of the annual observatory shutdown in 2021, but there is no record of any hardware or software changes that would have produced such a drastic CCF shape change. We confirm that this is not isolated to HD 185144 by computing the CCF FWHM time series for all stars that were observed during our precursor study in early 2021 (Figure 4). Nearly all stars see the same 25 m s-1 decrease in the average line width. We therefore find it necessary to define an additional RV era for data taken prior to 2021 August 31. We call this Run 0.5, electing to continue to use Run 1 to refer to data collected after this date and before the Contreras fire.
For CCF-based RV calculations, such as that of the NEID DRP, the RV zero point offsets resulting from these line profile changes will depend on additional factors such as projected rotational velocity () and effective temperature (), which affect the intrinsic RV information content of the stellar spectrum (Bouchy et al., 2001), on the CCF mask, which dictates which lines are used to calculate the RV, and on the systemic velocity (), which determines where these lines fall on the detector. Therefore, while the reference wavelength shifts are expected to be stable, the RV offset will vary from star to star.
@@
The above results were derived using late F-, G-, and early K-dwarfs, which make up the majority of the NETS sample. However, a chromatic (wavelength-dependent) zero point offset is hinted at by the FWHM time series for HD 95735 (Figure 4); this star was not used to calibrate the zero point offsets due to the presence of its long-period companion HD 95735 c ( d; Hurt et al., 2022). The FWHM for most stars decreases from Run 0.5 to Run 1 then holds relatively steady, though with a small increase, through Run 2. HD 95735, HD 170657, and HD 201091 deviate from this pattern. This can be explained by stellar activity in the case of HD 170657 and HD 201091, both of which exhibit increasing and highly variable chromospheric emission over this time frame. But activity levels remained stable for HD 95735 while the FWHM increases from Run 0.5 to Run 1 and increases again from Run 1 to Run 2. As the lone M-dwarf in the NETS target sample, HD 95735 has a CCF dominated by lines at longer wavelengths than those of the FGK-dwarfs that make up the rest of the sample. We check for any chromatic dependence by calculating the zero point offset as a function of wavelength, fitting the change in the mean stellar RV for each order over the wavelength range 380 nm – 790 nm (echelle orders 162–78). For the Run 0.5 to Run 1 zero point offset (Figure 5), we show that our data are insufficient to measure any significant chromatic variation, as the offset for each order is consistent with zero. For the Run 1 to Run 2 offset, however, we do observe an interesting chromatic signal (Figure 6). In particular, we find large offset variations from order to order in the wavelength region shorter than 530 nm, anchored by the ThAr calibration lamp. We attribute this to the new ThAr master wavelength solution. Further, the offset across all orders anchored by ThAr is m s-1 while the offset across the redder orders, anchored by the laser frequency comb (LFC), is m s-1. One should therefore exercise caution when using the mean offsets calculated for the full spectrum to inform analyses of M-dwarfs, which derive very little RV information content from the ThAr orders. Our results suggest redder stars should be insensitive to the Run 1 to Run 2 RV offset, though this needs to be confirmed via a similar analysis using a sample of M-dwarfs.

### v2

@@ 1 Introduction
The NEID Earth Twin Survey (NETS; A. F. Gupta et al., 2021) is a dedicated search for low-mass, long-period exoplanets with NEID (C. Schwab et al., 2016), an extreme precision radial velocity (EPRV) spectrograph on the WIYN 3.5 m telescope at Kitt Peak National Observatory.111The WIYN Observatory is a joint facility of the NSF’s National Optical-Infrared Astronomy Research Laboratory, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, and Princeton University. Using time allocated through a guaranteed time observations (GTO) program, supplemented in part by Penn State institutional access to NEID, we have been monitoring a carefully selected sample of 41 bright, RV-quiet stars since the commencement of this survey in September 2021. Data from the first few years of the survey have led to the confirmation of a low-mass planet (A. F. Gupta et al., 2025), the precise characterization of long-period stellar and planetary orbits (M. R. Giovinazzi et al., 2025), and the refinement of stellar activity mitigation methods (A. Burrows et al., 2024; J. C. Siegel et al., 2024; C. Gilbertson et al., 2024).
@@ 2.1 NETS Observing Strategy
NEID is a fiber-fed (S. Kanodia et al., 2023), environmentally-stabilized (P. Robertson et al., 2019) optical to near-infrared spectrograph. The instrument was designed to exceed an internal precision requirement of 27 cm s-1 based on a comprehensive, bottom-up error budget presented by S. Halverson et al. (2016). Tests performed during instrument commissioning and data from the NEID solar feed (A. S. J. Lin et al., 2022; E. B. Ford et al., 2024) demonstrate an achieved on-sky precision of cm s-1. In addition, thanks to its broad wavelength coverage of more than 380 – 930 nm, NEID observations cover a number of spectral features that are known to trace stellar magnetic activity, including the Ca II H&K lines, H, and the Ca II infrared triplet.
@@ 2.1.1 Nightly Observing Strategy
The target single-measurement precision for NETS observations follows that described by A. F. Gupta et al. (2021), which we summarize here. We first calculate how long it will take to reach a combined cm s-1. The photon noise precision, , is estimated for median observing conditions using the NEID exposure time calculator, which interpolates over a grid of exposure time, stellar effective temperature (), and apparent -mag generated following F. Bouchy et al. (2001). For the photon noise calculation, the projected rotational velocity, , was held fixed at 2 km s-1. The residual p-mode oscillation amplitude, , is estimated using the model outlined by W. J. Chaplin et al. (2019). We then compare this to the period corresponding to the peak of the oscillation power envelope, , and take the larger of these two values to be the nominal exposure time for each star. To avoid incurring additional uncertainty due to variations in the charge transfer inefficiency (CTI) of the NEID detector (C. H. Blake et al., 2017), individual exposures are set to trigger at a fixed signal-to-noise ratio (S/N) corresponding to this nominal exposure time. For the brightest stars in our sample, for which the calculated S/N exceeds the detector non-linearity threshold at any wavelength (60% well depth; S/N per 1-D extracted pixel), observations are split into multiple consecutive exposures with a total baseline (including inter-exposure readout) equal to the nominal exposure time. We also note that for all NETS observations, we simultaneously feed light from the etalon (M. K. Kreider et al., 2025) through the calibration fiber, except when the etalon is offline for maintenance. These measurements can be used to improve sampling of the nightly instrument drift.
@@
The 30 cm s-1 precision threshold was selected by A. F. Gupta et al. (2021) because it is on par with the internal NEID precision requirement and it can be achieved with reasonably short exposures ( minutes) for most NETS targets. For a subset of these stars, however, we expect to reach 30 cm s-1 with exposure sequences shorter than the -minute overheads associated with each NEID observation. To improve the duty cycle and increase the precision of observations for some of the brightest stars, we double the number of exposures per visit for the stars HD 4614A, HD 10476, HD 26965, HD 115617, HD 185144, and HD 201091, and we triple the number of exposures per visit for HD 10700. In each of these cases, the total nominal visit duration is still less than 10 minutes.
@@
While NEID commissioning was ongoing and prior to initiating observations for the full NETS sample in September 2021, we observed a limited set of stars from January 2021 through August 2021 to explore the precision capabilities of the spectrograph and the practicality of our intended nightly observing strategy within the constraints of the NEID Queue.222For details about the NEID Queue, see https://wiyn-queuemaster.kpno.noirlab.edu/manual/ For these precursor observations, we aimed to schedule two visits per night, separated by a minimum of two hours. This sampling was intended to mitigate the impact of correlated noise contributions from granulation and supergranulation (X. Dumusque et al., 2011), as these processes have characteristic timescales of hours to days. While these observations were executed as intended in most cases, we were sometimes left with unpaired visits as a result of changing observing conditions and unforeseen technical issues. In addition, while the measured change in RV within a night was significant for some stars (e.g., HD 86728, as shown in A. F. Gupta et al., 2025), paired measurements were typically consistent to within the photon noise uncertainties for most targets. With fewer than ten paired measurements per star during this pre-survey phase, the results of the test were not conclusive. In the interest of maintaining a more consistent observing strategy across all nights, we elected to proceed with just a single visit per night once the full survey began.
@@ 2.1.2 Seasonal Observing Strategy
For each star, we schedule P1 visits in the middle of each season at high cadence, set to execute as frequently as once every two nights. These observations, which typically span 1–2 months, allow us to densely sample the RV time series on timescales comparable to the rotation periods of our target stars. Rotationally-modulated stellar activity signals are expected to complicate exoplanet detection for even the quietest stars (J. Crass et al., 2021), but simultaneous measurement of RVs and activity indicators can be a powerful tool for disentangling stellar and planetary signals at the sub-m s-1 level. However, this requires that our data sets are sufficiently well sampled so that we can measure and model activity variations on the relevant timescales (V. Rajpaul et al., 2015; R. D. Haywood et al., 2018; R. D. Hall et al., 2018; A. F. Lanza et al., 2019; H. Yu et al., 2024; A. F. Gupta & M. Bedell, 2024). On their own, these densely-sampled observations would lead to poor phase sampling for longer period exoplanet signals, which can degrade the accuracy and precision of the extracted signal amplitude (J. Burt et al., 2018). To improve our phase sampling, we use our P2 allocation to schedule low cadence observations that span each star’s observing season. Finally, to reduce annual aliasing effects (which can be a significant obstacle to the detection of planet with periods close to one year), we schedule sets of 2–3 P0 visits as each star is rising or setting to bracket the observing season. By making use of the highest priority time for these observations, we ensure that the stars are observed within a few weeks of each end of the observing season when they are observable for no more than a couple hours each night. This counteracts the tendency for targets to drop out of the schedule during these limited observability windows. A representative version of the seasonal observing schedule for a typical star, with ideal uniform spacing between visits, is shown in the top panel of Figure 1. The schedule is supplemented by additional P3 and P4 time provided by Penn State; these observations follow the same nightly observing strategy, but no constraints are imposed on the observing cadence at these low priorities. We allow each star to be observed no more than once per night across all priority levels.
@@ 2.2 Current Status of NETS Observations
NETS observations formally commenced in September 2021, at the start of the 2021B observing semester. We show the achieved NETS observing schedule through August 2024 in Figure 1. In addition to the observations that follow the seasonal and nightly strategies outlined above, we also include supplemental data taken outside of this algorithmic scope and data from the NEID standard star program. The supplemental data adhere to the nightly NETS observing strategy (Section 2.1.1) and these include observations at lower priorities (P3 and P4) as well as observations scheduled on an ad hoc basis to monitor interesting signals or to compensate for weather losses and other scheduling setbacks. The target list for the NEID standard star program includes the following NETS targets: HD 4628, HD 9407, HD 10700, HD 127334, HD 185144, and HD 221354. The standard star program uses the same S/N trigger as NETS for all but HD 10700, for which a slightly lower threshold is used. For the analysis in this work, we combine data from both programs, as the RV uncertainty introduced by CTI variations is not expected to exceed 10 cm s-1 for observations at high S/N (R. T. Blackman et al., 2020). For more detailed analyses, one could consider separating the NETS and standard star data for HD 10700.
@@ 2.3 Data Products
This work uses NETS data as processed by version 1.4 of the NEID Data Reduction Pipeline (DRP)444Pipeline documentation can be found at https://neid.ipac.caltech.edu/docs/NEID-DRP/, which uses the cross-correlation function (CCF; A. Baranne et al., 1996) method with spectral type-dependent masks to calculate RVs. We retrieved the processed spectra from the NEID data archive.555https://neid.ipac.caltech.edu Data collected during Science Run 1 (2020 October 26 – 2022 June 16) and Science Run 2 (2022 October 18 – 2024 August 19) are treated as distinct RV eras, as the thermal cycling of the spectrograph in response to the Contreras fire is expected to have introduced an offset in the instrumental RV zero point666https://neid.ipac.caltech.edu/docs/NEID-DRP/rveras.html. We have also identified an apparent RV zero point offset in the middle of Science Run 1, the physical cause of which is yet unknown, so we consider data taken from 2020 October 26 – 2021 August 31 as a separate “Run 0.5”. We refer to these as Run 0.5, Run 1, and Run 2 in this work, and the offsets are described in detail in Section 3.3.
In addition to precise RVs, the DRP provides useful diagnostics such as the CCF full width at half maximum (FWHM) and bisector inverse slope (BIS) as well as telluric-corrected measurements of various indicators of stellar activity levels. These are calculated as the ratio of the flux in the specified activity-sensitive line core relative to a nearby continuum region, as described in the DRP documentation. Among these indicators is the Ca II H&K emission, though the values provided by the NEID DRP have not been calibrated to the standard Mount Wilson (MW) scale for the commonly used SHK metric (D. K. Duncan et al., 1991, also called the S-index or S-value). We follow the approach of C. Lovis et al. (2011) and J. Gomes da Silva et al. (2021) and compute the scaled SHK,MW values from the native DRP Ca II H&K activity indicator by performing a linear fit between the average Ca II H&K values from the NEID DRP and the mean S-values from D. K. Duncan et al. (1991) for the 27 FGK stars that appear in both samples. We find a conversion S Ca II H&K, and we expect this calibration to be valid for stars with spectral types and activity levels comparable to the input sample (FGK stars with Ca II H&K). We report both the DRP Ca II H&K values as well as the scaled SHK,MW for each star, with the exception of HD 95735 (GJ 411), the sole M-dwarf on the NETS target list. The time series and generalized Lomb-Scargle (GLS; M. Zechmeister & M. Kürster, 2009, 2018) periodograms for several activity-sensitive lines are shown in Figure 2 for a typical NETS target, and in the accompanying figure set for the remaining target stars. False alarm probabilities (FAP) are calculated following M. Zechmeister & M. Kürster (2009), accounting for the number of independently sampled frequencies.
@@ 3.2 Stellar companions
Several of the stars in our sample are members of binary (or higher multiplicity) systems. To determine which of these stellar companions are expected to induce measurable RV signals, we use binary_mc (M. Giovinazzi, 2025) to simulate possible orbits based on the positions of each star as given by the Gaia DR3 catalog (Gaia Collaboration et al., 2023). We calculate the maximum RV slope for each of 10 million random orbit draws for each system, and we consider the slope to be measurable in systems for which at least 10% of orbits have maximum slopes m s-1 yr-1. Four of these, HD 4614 A, HD 24496 A, HD 68017, and HD 201091 (61 Cyg A), were analyzed in a study of NETS stars with high astrometric accelerations by M. R. Giovinazzi et al. (2025) which combined astrometric measurements, archival RVs, and Run 1 NETS data to refine the binary orbits. We adopt and subtract the published orbits from the observed RVs for each of these stars. The orbital parameters are given in Table 2. An additional three stars, HD 179957, HD 186408, and HD 116442 have measurable RV signals due to their stellar companions but did not meet the criteria for thorough characterization in M. R. Giovinazzi et al. (2025). For these stars, we fit and subtract separate linear trends from NEID time series for each run.
Other stars in our sample with stellar-mass companions include HD 86728, which has a low-mass M-dwarf binary (J. E. Gizis et al., 2000; L. A. Hirsch et al., 2021) and HD 26965, which is part of a triple star system in which the orbit of the B and C system is well characterized (B. D. Mason et al., 2017) but that of the BC pair about the primary star is not. For these systems, we neither expect nor observe measurable RV signals as a consequence of some combination of long orbital periods, low companion masses, and highly inclined orbits.
@@ 3.3 RV Zero Point Offset
The extraordinary stability of EPRV spectrographs such as NEID is achieved through careful environmental control with extremely strict tolerances (e.g., G. Stefansson et al., 2016; P. Robertson et al., 2019; J. Crass et al., 2022). As a consequence, changes to the reference wavelength scale and average line profile resulting from even minor changes to the spectrograph or its calibration sources may introduce a measurable RV zero point offset. Our analysis of the line profiles of NETS observations reveals that two such changes occurred during the first three years of the survey. In Figure 3, we show the residual CCFs for all spectra of the star HD 185144, which was observed as part of both the NEID standard star program and NETS. The residual CCFs are calculated as the difference between the normalized CCF of each observation and the average normalized CCF over all observations. There are two distinct changes to the average line profile; the first occurs between 2021 August 7 and 2021 September 17, and the second occurs between 2022 June 10 and 2022 December 9. The latter change corresponds to the break between Run 1 to Run 2 and is explained by the derivation of a new master wavelength solution for the Thorium-Argon (ThAr) calibration lamp and by the thermal cycling of the instrument in response to the Contreras fire. The earlier break corresponds to dates of the annual observatory shutdown in 2021, but there is no record of any hardware or software changes that would have produced such a drastic CCF shape change. We confirm that this is not isolated to HD 185144 by computing the CCF FWHM time series for all stars that were observed during our precursor study in early 2021 (Figure 4). Nearly all stars see the same 25 m s-1 decrease in the average line width. We therefore find it necessary to define an additional RV era for data taken prior to 2021 August 31. We call this Run 0.5, electing to continue to use Run 1 to refer to data collected after this date and before the Contreras fire.
For CCF-based RV calculations, such as that of the NEID DRP, the RV zero point offsets resulting from these line profile changes will depend on additional factors such as projected rotational velocity () and effective temperature (), which affect the intrinsic RV information content of the stellar spectrum (F. Bouchy et al., 2001), on the CCF mask, which dictates which lines are used to calculate the RV, and on the systemic velocity (), which determines where these lines fall on the detector. Therefore, while the reference wavelength shifts are expected to be stable, the RV offset will vary from star to star.
@@
The above results were derived using late F-, G-, and early K-dwarfs, which make up the majority of the NETS sample. However, a chromatic (wavelength-dependent) zero point offset is hinted at by the FWHM time series for the lone M-dwarf in our sample, HD 95735 (Figure 4); this star was not used to calibrate the zero point offsets due to the presence of its long-period companion HD 95735 c ( d; S. A. Hurt et al., 2022). The FWHM for most stars decreases from Run 0.5 to Run 1 then holds relatively steady, though with a small increase, through Run 2. HD 95735, HD 170657, and HD 201091 deviate from this pattern. This can be explained by stellar activity in the case of HD 170657 and HD 201091, both of which exhibit increasing and highly variable chromospheric emission over this time frame. But activity levels remained stable for HD 95735 while the FWHM increases from Run 0.5 to Run 1 and increases again from Run 1 to Run 2. As an M-dwarf, HD 95735 has a CCF dominated by lines at longer wavelengths than those of the FGK-dwarfs that make up the rest of the sample. Using the same subset of stars as before, we check for any chromatic dependence by calculating the zero point offset as a function of wavelength, fitting the change in the mean stellar RV for each order over the wavelength range 380 nm – 790 nm (echelle orders 162–78). For the Run 0.5 to Run 1 zero point offset (Figure 5), we show that our data are insufficient to measure any significant chromatic variation, as the offset for each order is consistent with zero. For the Run 1 to Run 2 offset, however, we do observe an interesting chromatic signal (Figure 6). In particular, we find large offset variations from order to order in the wavelength region shorter than 530 nm, anchored by the ThAr calibration lamp. We attribute this to the new ThAr master wavelength solution. Further, the offset across all orders anchored by ThAr is m s-1 while the offset across the redder orders, anchored by the laser frequency comb (LFC), is m s-1. One should therefore exercise caution when using the mean offsets calculated for the full spectrum to inform analyses of M-dwarfs, which derive very little RV information content from the ThAr orders. Our results suggest redder stars should be insensitive to the Run 1 to Run 2 RV offset, though this needs to be confirmed via a similar analysis using a sample of M-dwarfs.